# SplatVoxel: History-Aware Novel View Streaming without Temporal Training

### v1

@@ 5.1 Feed-forward Novel View Synthesis
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
GT GS-LRM Ours
Figure 5: Reconstrution Close-ups. We show zoomed-in novel view reconstructions of GS-LRM and our model on dynamic scene datasets. Our model better handles occlusion boundaries with sharper detail.
Figure 6: Temporal Slice Comparison. We extract a horizonal line in the ground truth and generated videos to visualize temporal stability. Applying GS-LRM independently per frame is more prone to flickering, while our method achieves greater stability and is more robust to occlusion.
5.2 Novel View Streaming
We compare our method with baselines on the novel view streaming task on two dynamic scene datasets, Neural3DVideo [26] and LongVolumetricVideo [75]. We use 2 input views and 290 frames for the six sequences in Neural3DVideo, and 4 input views with 1500 frames for the two sequences in LongVolumetricVideo. For quantitative evaluation, we use a temporal metric to measure flicker in the rendered streaming videos in addition to the aforementioned per-frame novel view reconstruction metrics. Temporal flickering is measured using feature-space differences over time. Given the predicted frames $I_{t}$ and ground truth frames $I_{t}^{*}$, we extract deep features using a pretrained VGG network [55] $\phi(\cdot)$:
$\text{Flicker}_{t}=\big{|}D_{t}-D_{t}^{*}\big{|}$ (11)
where $D_{t}=\|\phi(I_{t})-\phi(I_{t-1})\|_{2}$ and $D_{t}^{*}=\|\phi(I_{t}^{*})-\phi(I_{t-1}^{*})\|_{2}$ represent the temporal feature differences for the predicted and ground truth frames at frame $t$, respectively. The final flicker score is computed by averaging $\text{Flicker}_{t}$ over the entire sequence. Lower values indicate smoother temporal consistency.
Compared to our proposed history-aware feed-forward method, other baseline methods can be classified into three categories: per-frame optimization, temporal optimization, and per-frame feed-forward reconstruction. As presented in Table 3, our method achieves the highest per-frame novel view synthesis performance across all three evaluation metrics on both datasets with the second-best flicker scores. Notably, our approach significantly enhances novel view synthesis results on the Neural3DVideo dataset, where only two input views are available, demonstrating its robustness in handling extremely sparse input settings. In comparison to the state-of-the-art per-frame feed-forward method GS-LRM, our approach qualitatively reconstructs geometry and appearance with greater accuracy and finer details (Fig. 4), and we show close-ups of the reconstruction in Fig. 5. Our method reduces flickering through history-aware modeling, visualized with temporal slices (Fig. 6). Although 4D Gaussian Splatting achieves a better flicker metric, the novel view quality is substantially worse compared to ours. Moreover, its time-consuming temporal optimization is impractical for our target application, which focuses on online processing of arbitrarily long sequences.
5.3 Ablation Studies
We show model ablations on DL3DV in Table 4. Naive mean averaging of Gaussian attributes in each voxel significantly degrades results. The absence of a coarse-to-fine setup limites the capacity of the output Splats, and without sparsification, we downsampled the grid to 0.75x its original resolution to avoid exceeding memory capacity for the voxel transformer, both leading to poorer reconstruction quality. We also experiment using a 3D CNN to downsample the voxel grid instead of our proposed coarse-to-fine sprase voxel grid, which is $\sim$1 dB worse than our final result and much slower. We also find splat features helpful, operating on splat parameters directly leads to $\sim$1 dB drop in PSNR.
We performed ablation studies on the Neural3DVideo dataset to evaluate our streaming reconstruction components, as shown in Table 5. Removing historical Gaussians fails to resolve temporal flicker. Disabling 3D warping led to abrupt keyframe changes and increased reconstruction errors. Eliminating error-aware fusion degraded reconstruction quality due to warping artifacts and other errors.
Table 5: Ablation Studies on Temporal Components.
@@ 6 Conclusion
In summary, we introduce a novel approach that leverages information from previous frames to enhance novel view streaming with sparse cameras. Our proposed SplatVoxel framework employs a hybrid representation by modeling scenes with Gaussian primitives and fusing them across frames using sparse voxel grids. This design enables training only on static scenes while supporting zero-shot inference on multi-view video streams via explicit 3D scene flow warping. As a result, our method achieves state-of-the-art performance on multiple multi-view video datasets under sparse camera setups and runs at interactive rates on a single H100 GPU.
While our approach marks an important step toward real-time novel view streaming, it does have limitations. Using pretrained 2D tracking models can be slow and prone to errors in challenging scenarios. Moreover, the current framework still performs a partial 3D reconstruction at each frame rather than incrementally updating previous reconstructions. Future work can explore strategies to leverage historical information without explicit tracking and incrementally refine prior reconstructions, potentially prioritizing regions with higher reconstruction errors. More efficient use of historical reconstructions could also improve the handling of high-resolution novel view streaming. These avenues promise to further enhance the efficiency and quality of history-aware novel view streaming.
@@ Acknowledgement.
We would like to express our gratitude to John Flynn, Kathryn Heal, Lynn Tsai, Srinivas Kaza, and Tooba Imtiaz for their helpful discussions, as well as Clément Godard, Jason Lawrence, Michael Broxton, Peter Hedman, Ricardo Martin-Brualla, and Ryan Overbeck for their valuable comments.
Supplementary Material
@@ Appendix A Implementation Details
Our framework is implemented in JAX [3] and trained on NVIDIA A100 GPUs. We evaluate the speed of both GS-LRM and our method on a single NVIDIA H100 GPU. We evaluate the speed of other optimization-based baselines, including 3DGS, 3DGStream, and 4DGS, on a single NVIDIA A100 GPU due to limitations with the H100 card. While running on an H100 can offer speed improvements, these baselines remain constrained by their inherent computational complexity and fail to achieve 1FPS.
We use the Adam optimizer [23] with an initial learning rate of 4e-4, applying cosine learning rate decay with linear warmup. The warmup period is set to 5000 training steps.
@@ A.1 Network Structure
Each transformer layer [66] consists of a self-attention mechanism and feed-forward networks, with all bias terms removed following GS-LRM [79]. Additionally, following LVSM [20], we incorporate QK-Norm [16] to stabilize training. For efficiency, we utilize the CUDNN-FlashAttention [9] implementation available in Jax.
The shallow multi-layer perceptrons $\mathrm{MLP}_{coarse}$ and $\mathrm{MLP}_{fine}$ share similar structures with the feed-forward network in a transformer layer. It includes two linear layers with a LayerNorm [1] and GeLU [15] activation function,and an additional residual connection.
@@ A.2 Hybrid Splat-Voxel Representation
Our SplatVoxel model is inspired by Lagrangian-Eulerian representations in physical simulation where we combine the gaussian splats (Lagrangian representation) and voxels (Eularian representation) to assist warping and fusion of history information.
We use the offical splatting-based rasterizer from 3D Gaussian Splatting [22] to render images from predicted Gaussian primitives. To ensure a fair comparison with GS-LRM on RealEstate10K[82], we set the spherical harmonics degree to 0 for Table 6. For models trained on the DL3DV dataset [37], the degree is set to 1 to enhance view-dependent effects.
Our voxel grid is aligned with the target camera, and is in the normalized device space (NDC), where the z-axis is linear in disparity. This space is helpful in presenting unbounded, in-the-wild scenes. The full fine volume size is set to $[64,H,W]$, where the $H,W$ corresponds to the rendered image height and width. The coarse volume size is set to $[32,H//8,W//8]$, meaning each coarse-level volume is subdivided into $[2,8,8]$ fine-level volumes.
We initialize the voxels by depositing the Gaussian primitives to the volume with a 3D distance-based kernels used in Material Point Method [58]:
@@
where $\mathbf{x}=(x,y,z)$ is the offset from the voxel center to the Gaussian primitive center, and $h$ is the grid size used to normalize the offset vector to the range $[-1,1]$. The function $K(x)$ is a one-dimensional cubic kernel function given by:
@@
A.3 3D Warping
@@
We use TAPIR [10] as our 2D tracking backbone, which takes query points and video frames as input and produces 2D tracking positions throughout the video with the corresponding occlusion mask, indicating whether a tracked point is visible in a given frame. In addition to the occlusion mask obtained from 2D tracking, the visibility mask $\gamma_{i}$ used in Eqn. 10 for triangulation also depends on depth-based visibility. Depth-based visibility is determined by comparing the projected depth of a given splat with the depth map obtained by rendering all splats to a given view. If the projected depth is smaller, the splat is considered unoccluded. This process utilizes alpha blending of splat depths to derive rendered depths from 3D Gaussian Splatting, a widely adopted technique that has been demonstrated to be effective [7].
@@
Assuming motion continuity between neighboring frames, we first estimate optical flow between each keyframe and its subsequent frame using RAFT [64] to identify regions of significant movement. We select regions within the input views where the optical flow norm exceeds a threshold of 0.2 pixels as areas of significant motion. Each splat is then projected onto the input views to determine whether its projected position falls within these significant motion regions and remains unoccluded based on the depth-based visibility verification aforementioned. If a splat exhibits significant motion in more than half of the views, it is prioritized for sampling as an anchor splat. During experiments, we found that uniformly sampling 512 points from the prioritized splats using the Farthest Point Sampling method  [46] is sufficient for effectively handling moving Gaussians in the keyframes.
The deformation of each splat is computed using Linear Blend Skinning (LBS) based on the deformation of its associated anchor splats. For each Gaussian, we employ a K-nearest neighbors approach with $k=4$ to identify the closest anchor splats. The position displacement $\delta\mathbf{p}\in\mathbb{R}^{3}$ is then determined by blending the deformations of the selected anchor points, denoted as $\mathbf{t}_{i}\in\mathbb{R}^{3}$.
@@
Error-aware Temporal Fusion
Static areas are identified during the Splat-to-voxel transfer process, where historical splats undergoing deformation mark their neighboring voxels as regions of motion.
Appendix B Two-stage training
Figure 7: Comparison of Multi-View and Voxel Transformers. Validation curves of our transformer variations on the DL3DV dataset. Compared to the Multi-view transformer which produces gaussian primitives aligned with the 2D input images, our Sparse Voxel Transformer is able to converge at a better result within very few training steps.
Our training procedure first trains the multi-view transformer portion, and then jointly trains the multi-view transformer with the voxel transformer. As shown in the training curves 7, our proposed voxel transformer efficiently reasons about 3D information, surpassing the performance of only using the multi-view transformer with the same number of steps.
Table 6: Reproduced GS-LRM Performance. Our reproduced GS-LRM achieves improved LPIPS and SSIM scores with a slight decrease in PSNR on the RealEstate10K dataset compared to the officially reported numbers.
@@ Appendix B Two-stage training
pixelNeRF [78] 20.43 0.589 0.550
GPNR [59] 24.11 0.793 0.255
Du et al. [12] 24.78 0.820 0.213
pixelSplat [5] 26.09 0.863 0.136
MVSplat [6] 26.39 0.869 0.128
GS-LRM (Official) [79] 28.10 0.892 0.114
@@
Ours 28.47 0.907 0.078
Table 7: Quantitative evaluation comparing our method with the best-performing baselines, 4DGS [73] and GS-LRM [79], across all sequences in LongVolumetricVideo [75] and Neural3DVideo [26]. The best and second-best results are highlighted for clarity.
@@
Appendix C Dataset
C.1 Static Scene Datasets
RealEstate10K [82]

### v2

@@ 6.2 Novel View Streaming
Multiview Transformer 2D Tracking 3D Warping Fuse-and-Refine Rendering
Time (s) 0.03 0.28 0.018 0.021 0.003
Table 7: Ablation on Temporal Components. Incorporating historical information and our proposed mechanisms for leveraging it are crucial for achieving consistent and high-quality novel-view streaming results.
@@
Table 8: Comparison Between Two-stage and Joint training. Jointly training GS-LRM and our method achieves better performance than the sequential two-stage strategy that pre-trains GS-LRM.
[…]
GS-LRM 28.59 0.925 0.063
Ours (Two-Stage) 30.34 0.934 0.054
Ours (Joint) 30.61 0.935 0.052
6.3 Ablation Studies
Ablation on Sparse Voxel Transformer.
We show model ablations on DL3DV in Table 2. Naive non-learning fusion of Gaussian attributes in each voxel significantly degrades results. The absence of a coarse-to-fine setup limits the capacity of the output Splats, and without sparsification, we downsampled the grid to 0.75x its original resolution to avoid exceeding memory capacity for the voxel transformer, both leading to poorer reconstruction quality. We also experiment with using a 3D CNN to downsample and upsample the voxel grid instead of our designed coarse-to-fine voxel hierarchy. This approach yields results approximately 1 dB worse than our final model and is significantly slower. We also find splat features helpful, operating on splat parameters directly leads to $\sim$1 dB drop in PSNR.
Ablation on Temporal Components.
We conducted ablation studies on the Neural3DVideo dataset to evaluate our streaming reconstruction components, as shown in Table 8. Removing historical splats for fusion failed to resolve temporal flicker. Disabling 3D warping caused abrupt keyframe changes and higher reconstruction errors. Removing error-aware fusion degraded quality due to warping artifacts and other errors.
Ablation on Training Schemes.
Our framework supports both the two-stage training strategy described in Sec. 4.3 and joint training of the multi-view and voxel Transformers from scratch. As shown in Table 8, the joint training scheme can achieve better performance than the two-stage approach.
@@ 7 Conclusion
In summary, we introduce a novel Fuse-and-Refine module that can efficiently consolidate Gaussian primitives from multiple sources into a canonical and consistent representation. It enhances existing feed-forward models, and the design of the module naturally allows for reusing primitive information from past frames for better temporal coherence in streaming setups. Extensive evaluations show that our method achieves state-of-the-art performance on both static and dynamic datasets under sparse camera setups, delivering sharper reconstructions and reduced temporal flicker compared to previous feed-forward and optimization-based methods, all at interactive frame rates on a single GPU.
While the results of our approach are promising, it does have limitations. First, it depends on reasonably good initial primitives. For instance, in large-baseline setups such as the LongVolumetricVideo dataset, even the current best feed-forward model, GS-LRM, experiences a performance drop. As a result, our method yields limited improvements based on its input in such challenging scenarios. A potential solution is to train on a larger-scale dataset that covers a broader range of baseline configurations. Second, temporal artifacts still persist, particularly in regions with dynamic objects. The observed artifacts primarily arise from our current simple 3D warping strategy. Lifting 2D tracking into 3D from sparse views remains inherently error-prone and presents a persistent challenge in this ongoing field. A promising future direction is to investigate approaches that avoid explicit tracking or adopt adaptive online reconstruction strategies, potentially prioritizing regions with higher reconstruction errors.
@@ Acknowledgement.
We would like to express our gratitude to John Flynn, Kathryn Heal, Lynn Tsai, Srinivas Kaza, Qin Han, and Tooba Imtiaz for their helpful discussions, as well as Clément Godard, Jason Lawrence, Michael Broxton, Peter Hedman, Ricardo Martin-Brualla, and Ryan Overbeck for their valuable comments.
@@ Appendix A Implementation Details
Our framework is implemented in JAX jax2018github and trained on NVIDIA A100 GPUs. We use the Adam optimizer kingma2014adam with an initial learning rate of 4e-4, applying cosine learning rate decay with linear warmup. The warmup period is set to 5000 training steps.
@@ A.1 Network Structure
Each transformer layer vaswani2017transformer consists of a self-attention mechanism and feed-forward networks, with all bias terms removed following GS-LRM zhang2025gslrm. Additionally, following LVSM jin2024lvsm, we incorporate QK-Norm henry2020qknormquery to stabilize training. For efficiency, we utilize the CUDNN-FlashAttention dao2022flashattention implementation available in Jax.
The shallow multi-layer perceptrons $\mathrm{MLP}_{coarse}$ and $\mathrm{MLP}_{fine}$ share similar structures with the feed-forward network in a transformer layer. It includes two linear layers with a LayerNorm ba2016layernorm and GeLU hendrycks2016gelu activation function,and an additional residual connection.
@@ A.2 Hybrid Splat-Voxel Representation
Our SplatVoxel model is inspired by Lagrangian-Eulerian representations in physical simulation where we combine the Gaussian splats (Lagrangian representation) and voxels (Eulerian representation) to assist warping and fusion of history information.
We use the official splatting-based rasterizer from 3D Gaussian Splatting kerbl20233dgs to render images from predicted Gaussian primitives. To ensure a fair comparison with GS-LRM on RealEstate10Kzhou2018rel10k, we set the spherical harmonics degree to 0 for Table 9. For models trained on the DL3DV dataset ling2024dl3dv, the degree is set to 1 to enhance view-dependent effects.
Our voxel grid is aligned with the one target camera, and is in the normalized device space (NDC), where the z-axis is linear in disparity. This space is helpful in presenting unbounded, in-the-wild scenes. The full fine volume size is set to $[64,H,W]$, where the $H,W$ corresponds to the rendered image height and width. The coarse volume size is set to $[32,H//8,W//8]$, meaning each coarse-level volume is subdivided into $[2,8,8]$ fine-level volumes. The cost volume can also be constructed from any meaningful reference view, such as the average of the input cameras, to produce 3D Gaussian Splatting (3DGS) for free-viewpoint novel view synthesis. Moreover, the application of our sparse voxel transformer is not limited to cost volumes. It naturally generalizes to other voxel structures, such as octrees, making it applicable to broader scenarios like SLAM.
We initialize the voxels by depositing the Gaussian primitives to the volume with a 3D distance-based kernels used in Material Point Method steffen2008MPManalysis:
@@
where $\mathbf{x}=(x,y,z)$ is the offset from the voxel center to the Gaussian primitive center, and $h$ is the grid size used to normalize the offset vector to the range [1,1][-1,1]. The function $K(x)$ is a one-dimensional cubic kernel function given by:
@@
Figure 5: History-aware Streaming System. Our method can generalize to dynamic scenes during inference while trained only on static scenes. The hybrid Splat-Voxel model first extracts input image features using a multi-view transformer, which outputs pixel-aligned Gaussian splats for each input image with associated features. The splat features are then deposited onto a coarse-to-fine voxel grid using the decoded position, and a secondary sparse voxel transformer processes the grid features to output final Gaussian parameters. To merge history, we compute the triangulated scene flow from the input views and perform keypoint guided deformations. These deformed splats can be treated identically to the input-aligned splats, and be similarly deposited into voxel grid to merge the previous state with the current state.
A.3 History-aware Streaming Reconstruction
@@
For solving the close closest point to the $N$ camera rays originating from the 2D projections $\mathbf{p}_{t}^{i}$, which can be formulated as the following least-squares optimization problem:
where $\mathbf{o}_{i}$ and $\mathbf{d}_{i}$ denote the origin and direction of the $i$-th camera ray, respectively, and $\lambda_{i}$ is a scalar representing the depth along the ray. This least-squares problems can be converted into solving an linear equation as
We use TAPIR doersch2023tapir as our 2D tracking backbone, which takes query points and video frames as input and produces 2D tracking positions throughout the video with the corresponding occlusion mask, indicating whether a tracked point is visible in a given frame. In addition to the occlusion mask obtained from 2D tracking, the visibility mask $\gamma_{i}$ used in Eqn. 14 for triangulation also depends on depth-based visibility. Depth-based visibility is determined by comparing the projected depth of a given splat with the depth map obtained by rendering all splats to a given view. If the projected depth is smaller, the splat is considered unoccluded. This process utilizes alpha blending of splat depths to derive rendered depths from 3D Gaussian Splatting, a widely adopted technique that has been demonstrated to be effective chung2024depthreg.
@@
Assuming motion continuity between neighboring frames, we first estimate optical flow between each keyframe and its subsequent frame using RAFT teed2020raft to identify regions of significant movement. We select regions within the input views where the optical flow norm exceeds a threshold of 0.2 pixels as areas of significant motion. Each splat is then projected onto the input views to determine whether its projected position falls within these significant motion regions and remains unoccluded based on the depth-based visibility verification aforementioned. If a splat exhibits significant motion in more than half of the views, it is prioritized for sampling as an anchor splat. During experiments, we found that uniformly sampling 512 points from the prioritized splats using the Farthest Point Sampling method  qi2017pointnet++ is sufficient for effectively handling moving Gaussians in the keyframes.
To warp splats from the previous frame, we first compute deformations on the anchor splats using triangulation, and then propagate the motion to all primitives through Linear Blend Skinning (LBS). For each Gaussian, we employ a K-nearest neighbors approach with $k=4$ to identify the closest anchor splats. The position displacement $\delta\mathbf{p}\in\mathbb{R}^{3}$ is then determined by blending the deformations of the selected anchor points, denoted as $\mathbf{t}_{i}\in\mathbb{R}^{3}$.
@@
For efficiency, we only maintain primitives from a set of past keyframes, and the contribution of these primitives are smoothly adjusted as older keyframes are replaced by newer ones to ensure seamless temporal transitions. Due to time constraints from querying the 2D tracking model, in practice we use two keyframes placed 5 frames apart, each lasting for 10 frames.
Error-aware Streaming Fusion
To further mitigate artifacts resulting from warping inaccuracies of historical primitives (e.g. warping error, appearance variation, or new objects), we use an adaptive fusion weighting mechanism during splat fusion based on the per-pixel error between the past and current frame splats when rendered to the input views. In more detail, each historical splat is projected onto the input views to assess whether it produces a higher rendering error and appears in front of the rendered depth of the current frame splats. If a historical splat results in increased error in more than half of the input views, we set its fusion weight to zero to suppress unreliable contributions. Concurrently, current frame splats predicted from the multi-view transformer that increase rendering error are assigned zero weight. To enhance temporal continuity, we retain 50% of the voxel attributes from the past keyframe in static areas. Static areas are identified during the Splat-to-voxel transfer process, where historical splats undergoing deformation mark their neighboring voxels as regions of motion.
Fig. 5 shows the system implementation of our history-aware streaming reconstruction. The streaming fusion and refinement process enables our model to significantly reduce temporal flickering and improve reconstruction quality, as demonstrated in Fig. 6.
Table 9: Reproduced GS-LRM Performance. Our reproduced GS-LRM achieves improved LPIPS and SSIM scores with a slight decrease in PSNR on the RealEstate10K dataset compared to the officially reported numbers. We also include comparisons with other feed-forward novel view synthesis methods in this table.
@@
pixelNeRF yu2021pixelnerf 20.43 0.589 0.550
GPNR suhail2022gpnr 24.11 0.793 0.255
Du et al. du2023learning 24.78 0.820 0.213
GS-LRM (Official) zhang2025gslrm 28.10 0.892 0.114
@@
GS-LRM (Reproduced) + Ours 28.47 0.907 0.078
Table 10: Inference Time on RealEstate10K. We compare the inference time of our method with other feed-forward Gaussian Splatting approaches. All runtimes are measured on an NVIDIA A100 GPU. The slight runtime overhead of our method relative to GS-LRM can be mitigated by reducing the number of multi-view transformer layers. As reported in Table 2, our method improves PSNR by over 2 dB while maintaining comparable model size and inference time.
[…]
pixelSplat charatan2024pixelsplat 26.09 0.863 0.136 0.104
MVSplat chen2024mvsplat 26.39 0.869 0.128 0.044
GS-LRM zhang2025gslrm 28.10 0.892 0.114 0.041
Ours 28.47 0.907 0.078 0.067
Table 11: Quantitative Comparison with 4DGS and GS-LRM. Evaluation across all sequences in LongVolumetricVideo xu2024longvolcap and Neural3DVideo li2022neural3DV. The best and second-best results are highlighted.
@@
Table 12: Method Comparison. Unlike existing baselines, our framework uniquely supports the three key requirements of the novel view streaming task targeted in this paper: fast reconstruction with interactive rate , sparse view inputs, and history awareness. In particular, our streaming approach only uses past frames, in constrast to DeformableGS yang2024deformable and DeformableNerf park2021nerfies that optimize over $\textit{all}^{*}$ frames. The history-awareness in our model mitigates temporal flickering and occlusion artifacts, and is also able to support unlimited sequence lengths.
Method  Interactive Speed  Sparse-view Input  History Awareness
NeRF mildenhall2021nerf $\times$ $\times$ $\times$
InstantNGP muller2022instantngp $\times$ $\times$ $\times$
3DGS kerbl20233dgs $\times$ $\times$ $\times$
DeformableGS yang2024deformable $\times$ $\checkmark^{*}$
DeformableNerf park2021nerfies $\times$ $\checkmark^{*}$
PixelSplat charatan2024pixelsplat $\times$
GS-LRM zhang2025gslrm $\times$
Quark flynn2024quark $\times$
Ours
Appendix B Dataset
B.1 Static Scene Datasets
RealEstate10K zhou2018rel10k